OpenStudy (anonymous):

given a polynomial has roots 2i + 3 and 2i -3 find p(x)

OpenStudy (anonymous):

is it really \(2i+3\) and \(2i-3\) ? very tricky if it is

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets write both of these numbers in standard form first they are \(3+2i\) and \(-3+2i\)

OpenStudy (anonymous):

if \(3+2i\) is a zero of a polynomial with real coefficients, then so is its conjugate \(3-2i\) now we can find a quadratic with those zeros

OpenStudy (anonymous):

the hard way is to write in factored form as \((x-(3+2i))(x-3(3-2i))\) and multiply out the easier way is to work backwards and put \[x=3+2i\\ x-3=2i\\ (x-3)^2=-4\\ x^2-6x+9=-4\\ x^2-6x+13=0\] and that is your quadratic with the two zeros \(3\pm2i\)

OpenStudy (anonymous):

the real easy way is to memorize that if \(a+bi\) is a zero of a quadratic, then the quadratic is \[x^2-2ax+(a^2+b^2)\] but that requires memorizing it in you case you would get \[x^2-2\times 3x+(3^2+2^2)=x^2-6x+13\] as before

OpenStudy (anonymous):

now we have to repeat the process with \(-3+2i\) this time we get \[x^2+6x+14\]

OpenStudy (anonymous):

and finally your last job is to multiply \[(x^2-6x+14)(x^2+6x+14)\]

OpenStudy (anonymous):

the answer i got was way out of pocket